OpenStudy (anonymous):

9^15-169 TIMES ABC = 95648 what is abc

thomaster (thomaster):

9^15 -169 * ABC = 95648 9^15 -169 = 95648 / ABC ABC = 95648 / 9^15 -169 ABC = 4.645*10^-10 Something like this?